4. a) Draw the perspective (dashes and wedges) structure of an androstane analog comprised of acarbonyl group at C-3, a carboxylic acid group at C-17 and a double bond between C-5 and C-6; whatkind of...

4.

a) Draw the perspective (dashes and wedges) structure of an androstane analog comprised of acarbonyl group at C-3, a carboxylic acid group at C-17 and a double bond between C-5 and C-6; whatkind of lipid is this?
b) Draw the perspective structure of the products of its reaction with 1) bromine followed by 2) water;assign R and S to all chiral centers. (Also consider whether these reagents react with the carbonyl and and/or the carboxylic acid group/s; can an HVZ reaction take place here?)
c) Draw the most stable chair conformation of each of the product(s)
Answer #1

a) linoleic acid UNSATURATED FAT (Linoleic Acid) Cis double bonds H-O-C-C-C-C-C-C=C-C-C= C-C-C-C-C-C-C-C-C-H TRANS FAT (trans-Linoleic Acid) Tb) reaction of linoleic with bromine

Br Br 2Brz (CH27CO2H CH3(CH2)4 Br CHa(CH2)4 Br Br Zn Br HCl (CH2COEt (CH2hCOH EtOH CH3(CH2)4 Br NaOH CH3(CH2)4

c)